# Scaling Analysis and Renormalization Group on the Mobility Edge in the Quantum Random Energy Model

### v1

@@ I Introduction
Since the early days of statistical mechanics, physicists have developed a very convenient way to count: zero, one, two, infinity. While sometimes all the numbers in between are easy to skip, and the thermodynamic limit can be accessed more or less straightforwardly, in many other situations finite-size corrections are very strong, resulting from interesting physical processes taking place on different scales. This is particularly true in the context of quantum localization transitions, arising from the competition between tunneling and disorder [1, 2]. Already in the single-particle setting, extensive numerical simulations have been required to extract reliable estimates of the critical exponents [3, 4, 5, 2]. Upon introducing interactions, the very existence of a stable many-body localized (MBL) phase has been debated for over 40 years [6, 7, 8, 9, 10, 11, 12, *Serbyn2013Local, 14, 15, 16], and only recently rigorous proofs have settled some aspects of the debate [17, *Imbrie2016Many, 19]. A great deal of controversy on the topic is due to the presence of severe finite-size effects [20, 21, 22, 23, 24, 25], masking the thermodynamic limit in numerical studies, and possibly giving rise to intermediate regimes [26, 27, 28]. In turn, analytical control for intermediate system sizes is poor because a reliable scaling theory of localization [29] in the presence of interactions is lacking, as the relevant—in the renormalization group (RG) sense—fields driving the transition have to be identified with confidence yet [30, 31, 32, 33, 34, 35, 36, 37].
Most of the analytical understanding of localization transitions is based on perturbative expansions. In the many-body case, the expansions are performed for effective single-particle models on Fock space graphs, under the assumption that loops have a negligible effect [8, 38]. For this reason, the Anderson model on expander graphs—such as random regular graphs (RRG)—has been argued to capture some of the phenomenology of the many-body case, and it has been studied extensively [39, 40, 5, 41, 42, 43, 44, 45, 46]. Recently, however, a RG approach that makes use of modern numerical techniques has shed new light on the problem [47]. A two-parameter scaling (2PS) hypothesis has been shown to describe the emergence of the localized phase, which corresponds to a line of fixed points, reminiscent of a Berezinskii-Kosterlitz-Thouless (BKT) scenario [48, *Berezinskii1972Destruction, 50]. This is in contrast with the transition in finite dimensions, where the fixed points of the transition, of the localized phase and of the ergodic phase are all described by a one-parameter scaling (1PS) RG flow [29, 51].
@@ II Model
The QREM is the quantum version of the random energy model, introduced by Derrida as an exactly solvable model arising from the limit $p\rightarrow\infty$ of $p$-spin glasses [57, *Derrida1981REM]. Its classical version ($\Gamma=0$) is typically considered as the simplest statistical mechanics model displaying a glassy phase below the critical temperature $T_{\mathrm{glass}}=1/2\sqrt{\log 2}$ [59]. Reintroducing quantum fluctuations, the equilibrium phase diagram of the model displays also a quantum paramagnetic phase, on top of the classical paramagnetic and glassy phases [60, 61, *Manai2023Spectral].
@@ III Spectral observables
The first is the spectral gap ratio, or $r$-parameter [9], $r=\mathbb{E}[\min(\Delta E_{n},\Delta E_{n+1})/\max(\Delta E_{n},\Delta E_{n+1})]$, where $\Delta E_{n}=E_{n+1}-E_{n}$ is the $n$-th energy gap in a narrow energy window and $\mathbb{E}$ denotes the average both over the window and the disorder. Notice that the use of such a window in the ergodic regime would define the microcanonical ensemble. The value $r_{\mathrm{P}}=2\ln 2-1\simeq 0.386$ is the Poisson value for integrable systems, while $r_{\mathrm{WD}}\simeq 0.5307$ is the Wigner-Dyson value for the Gaussian Orthogonal Ensamble (GOE), corresponding to chaotic systems with time-reversal symmetry [68]. For convenience, we rescale the average gap ratio defining $\phi=(r-r_{\mathrm{P}})/(r_{\mathrm{WD}}-r_{\mathrm{P}})$: $\phi=0$ corresponds to a non-ergodic behavior, while $\phi=1$ to ergodicity.
@@ IV Scaling theory of localization
In general, the beta function depends on all the relevant and irrelevant observables present in the theory. However, once the scaling regime is reached, the dependence on the irrelevant operators disappears, leaving only relevant observables to describe the proximity of the fixed points of the RG flow. If the critical regime is described by just a single relevant observable, i.e. there is one-parameter scaling (1PS), $\beta=\beta(A)$ holds. This happens, for instance, in the Anderson model in finite-dimensional Euclidean space [29, 51], where the localized point, the ergodic point and the transition point are all isolated fixed points. In the case in which the critical point collides with a line of fixed points, more parameters are needed to address the critical behavior: for instance, in the BKT [48, *Berezinskii1972Destruction, 50] transition or in the Anderson model in infinite dimensions [47] (and probably in interacting systems as well [55]) there are one relevant and one marginal operator. As a consequence, two equations are needed to describe the RG flow in the neighborhood of some critical points. In this case, the general equations describing two-parameter scaling (2PS) can be written using $\beta$ itself as a local coordinate:
@@ V Numerical results
Figure 2: (a) System-size dependence of the fractal dimension $D$ at the center of the spectrum (i.e. at infinite temperature). Dots are the numerical data, obtained by taking the finite-difference derivative of the participation entropy, and solid lines are the interpolation of the data, used for computing the beta function. For any $W$, the curves reach $D=1$ for sufficiently large system size. At intermediate $L$, the fractal dimension becomes $D>1$, signaling that the eigenfunction support set grows faster than the system size. (b) Beta function at $E=0$, with direction of the RG flow indicated by the arrows. The red dots in the main plot are the zeros of the beta function at $D>1$, and we denote with $D_{B}$ the corresponding value of $D$. In the inset, we report the dependence of $D_{B}$ on $W$: a linear behavior is observed, suggesting the persistence of this effect for large sizes. (c) Pictorial representation of the two possible scenarios for the RG flow at the ergodic fixed point, one-parameter (1PS) or two-parameter (2PS) scaling. The red dashed line is a sketch for the 1PS asymptotic curve.
Numerical data qualitatively agree with the picture coming from perturbation theory [52, 53]. The fractal dimension at the center of the spectrum, $E=0$, is reported as a function of system size in Fig. 2(a), and the corresponding beta function in Fig. 2(b): all curves flow to the ergodic fixed point for any value of $W$, as long as one keeps $W=O(1)$. The same conclusion can be reached from the $r$-parameter data, that we report in the Supplement [71]. We anticipate that one can observe a transition for $E=0$, provided one rescales the disorder as already pointed out in Refs. [52, 53]. We will explore this transition in the next section.
@@ VI Rescaling the disorder
We plot the beta function of the fractal dimension, parametrized by the rescaled disorder $\overline{W}$, in Fig. 3. While bigger system sizes are needed to draw definitive conclusions on the nature of the transition, it appears plausible that it is described by a 2PS theory, with a line of fixed points on the negative $\beta$ axis. The situation appears similar to the Anderson transition on expander graphs [47], and potentially to the random-field XXZ spin chain [55]. The approach to the ergodic fixed point follows instead a 1PS curve, as the beta function curves collapse on one another.
The turning points $\phi_{A}$ of the RG dynamics, i.e. the minima of $\phi$, decrease linearly to $\phi_{c}=0$ at the finite value $\overline{W}_{c}\simeq 1.95$, when plotted as a function of $\overline{W}$ (see inset of Fig. 3(a)). Note that, with our definitions of the disorder strength, the critical disorder on a Bethe lattice would be $\overline{W}_{c}^{\mathrm{BL}}=2/\sqrt{\pi}\simeq 1.12$ [43]. It is possible that the value in the QREM turns out to be larger because of the presence of loops in the Fock-space graph, which favor resonances. Here loops are particularly important because the number of dimensions of the Fock space is increased while keeping the side of the hypercubic lattice fixed to 1. The Bethe lattice, instead, is reached by keeping the number of spins fixed and increasing the spin representation: this is equivalent to considering the Anderson model in finite $d$-dimensional space [51] and taking the limit $d\rightarrow\infty$. A possible way to interpolate the two situations would be to consider, instead of spins-1/2, an arbitrary spin $S$ which generates an $L$-dimensional hypercube with $(2S+1)^{L}$ vertices of size $12S+$ 1 in each direction. We conjecture that the critical value $\overline{W}_{c}(S)\to\overline{W}_{c}^{\mathrm{BL}}$ as $S\to\infty$.
@@ VII On the mobility edge
Figure 4: (a) Beta function (main) and the corresponding fractal dimension (inset) away from the center of the spectrum, at $E=E_{\mathrm{max}}/4$ with $E_{\mathrm{max}}$ largest energy eigenvalue. The approach to $D=1$ suffers from less severe finite-size effects, and $D$ only slightly exceeds 1. Moreover, at large $W$ the curves $D(L)$ approach $D=0$, signaling the presence of the localized phase. (b) (Main) Determination of the critical value $W_{c}$ at energy $E=E_{\mathrm{max}}/4$. In the plot, $\phi_{A}$ is such that $\beta(\phi_{A})=0$, and it extrapolates to $\phi_{A}=0$ for $W_{c}\simeq 20$. At $W=W_{c}$, quantities as the fractal dimension $D$ (inset) vanish as $1/L^{2}$, a behavior typical of expander graphs [45, 47].
@@ VIII Discussion
The scaling theory of the localization transition in the QREM has been studied in the present work by reconstructing numerically the beta function of spectral observables. In the “natural” scaling of the disorder variance, which makes the model a mean-field version of MBL models, our analysis confirms the absence of transition at the center of the spectrum and opens two possible scenarios for the approach to the ergodic fixed point: either a one- or a two-parameter scaling. On the other hand, at finite energy density a localization transition is described by a two-parameter scaling theory similarly to the Anderson model on expander graphs. A transition to a localized phase emerges also at the center of the spectrum if one suitably rescales the variance of the disorder: the transition appears to be described by a two-parameter scaling theory even in this case, with the same critical exponent of the Anderson model on the RRG and possibly of the XXZ spin chain. Our results show that the universality class of the localization transition in the QREM is the same of the Anderson model on any random graph and, remarkably, does not change by rescaling the disorder. This is a consequence of the fact that the RG is blind to the scaling of the microscopic variables.
@@ Acknowledgements.
We acknowledge discussion with C. L. Baldwin, I. V. Gornyi, D. A. Huse, C. R. Laumann, A. D. Mirlin and T. Scoquart, and we thank B. L. Altshuler and V. E. Kravtsov for collaborations on related topics. Numerical simulations were performed using the libraries SciPy [77], PETSc [78] and SLEPc [79] on the MPIPKS HPC cluster. This work was partly funded by the European Union – NextGenerationEU under the project NRRP “National Centre for HPC, Big Data and Quantum Computing (HPC)” CN00000013 (CUP D43C22001240001) [MUR Decree n. 341–15/03/2022] – Cascade Call launched by SPOKE 10 POLIMI: “CQEB” project. The work of JN was funded by the European Union–NextGenerationEU under the project NRRP Project “National Quantum Science and Technology Institute" — NQSTI, Award Number: PE00000023, Concession Decree No. 1564 of 11.10.2022 adopted by the Italian Ministry of Research, CUP J97G22000390007.

### v2

@@ I Introduction
Since the early days of statistical mechanics, physicists have been interested in understanding how the collective behavior of many degrees of freedom leads to the macroscopically observed phenomena. While sometimes the thermodynamic limit can be accessed more or less straightforwardly, in many other situations finite-size corrections are very strong, resulting from interesting and non-trivial physical processes taking place on different scales. This is particularly true in the context of quantum localization transitions, arising from the competition between tunneling and disorder [1, 2]. Already in the single-particle setting, extensive numerical simulations have been required to extract reliable estimates of the critical exponents [3, 2, 4, 5]. Upon introducing interactions, the very existence of a stable many-body localized (MBL) phase has been debated for over 40 years [6, 7, 8, 9, 10, 11, 12, 13, *Serbyn2013Local, 15, 16, 17, 18], and only recently rigorous proofs have settled some aspects of the debate [19, *Imbrie2016Many, 21]. A great deal of controversy on the topic is due to the presence of severe finite-size effects [22, 23, 24, 25, 26, 27], masking the thermodynamic limit in numerical studies, and possibly giving rise to intermediate regimes [28, 29, 30]. In turn, analytical control for intermediate system sizes is poor because a reliable scaling theory of localization [31] in the presence of interactions is lacking, as the relevant—in the renormalization group (RG) sense—fields driving the transition have to be identified with confidence yet [32, 33, 34, 35, 36, 37, 38, 39].
Most of the analytical understanding of localization transitions is based on perturbative expansions. In the many-body case, the expansions are performed for effective single-particle models on Fock space graphs, under the assumption that loops have a negligible effect [9, 40]. For this reason, the Anderson model on expander graphs—such as random regular graphs (RRG)—has been argued to capture some of the phenomenology of the many-body case, and it has been studied extensively [41, 42, 5, 43, 44, 45, 46, 47, 48]. Recently, however, a RG approach that makes use of modern numerical techniques has shed new light on the problem [49]. A two-parameter scaling (2PS) hypothesis has been shown to describe the emergence of the localized phase, which corresponds to a line of fixed points, reminiscent of a Berezinskii-Kosterlitz-Thouless (BKT) scenario [50, *Berezinskii1972Destruction, 52]. This is in contrast with the transition in finite dimensions, where the fixed points of the transition, of the localized phase and of the ergodic phase are all described by a one-parameter scaling (1PS) RG flow [31, 53].
@@ II Model
The QREM is the quantum version of the random energy model, introduced by Derrida as an exactly solvable model arising from the limit $p\rightarrow\infty$ of $p$-spin glasses [59, *Derrida1981REM]. Its classical version ($\Gamma=0$) is typically considered as the simplest statistical mechanics model displaying a glassy phase below the critical temperature $T_{\mathrm{glass}}=1/2\sqrt{\log 2}$ [61]. Reintroducing quantum fluctuations, the equilibrium phase diagram of the model displays also a quantum paramagnetic phase, on top of the classical paramagnetic and glassy phases [62, 63, *Manai2023Spectral].
@@ III Spectral observables
The first is the spectral gap ratio, or $r$-parameter [10], $r=\mathbb{E}[\min(\Delta E_{n},\Delta E_{n+1})/\max(\Delta E_{n},\Delta E_{n+1})]$, where $\Delta E_{n}=E_{n+1}-E_{n}$ is the $n$-th energy gap in a narrow energy window and $\mathbb{E}$ denotes the average both over the window and the disorder. Notice that the use of such a window in the ergodic regime would define the microcanonical ensemble. The value $r_{\mathrm{P}}=2\ln 2-1\simeq 0.386$ is the Poisson value for integrable systems, while $r_{\mathrm{WD}}\simeq 0.5307$ is the Wigner-Dyson value for the Gaussian Orthogonal Ensemble (GOE), corresponding to chaotic systems with time-reversal symmetry [70]. For convenience, we rescale the average gap ratio defining $\phi=(r-r_{\mathrm{P}})/(r_{\mathrm{WD}}-r_{\mathrm{P}})$: $\phi=0$ corresponds to a non-ergodic behavior, while $\phi=1$ to ergodicity.
@@ IV Scaling theory of localization
In general, the beta function depends on all the relevant and irrelevant observables present in the theory. However, once the scaling regime is reached, the dependence on the irrelevant operators disappears, leaving only relevant observables to describe the proximity of the fixed points of the RG flow. If the critical regime is described by just a single relevant observable, i.e. there is one-parameter scaling (1PS), $\beta=\beta(A)$ holds. This happens, for instance, in the Anderson model in finite-dimensional Euclidean space [31, 53], where the localized point, the ergodic point and the transition point are all isolated fixed points. In the case in which the critical point collides with a line of fixed points, more parameters are needed to address the critical behavior: for instance, in the BKT [50, *Berezinskii1972Destruction, 52] transition or in the Anderson model in infinite dimensions [49] (and probably in interacting systems as well [57]) there are one relevant and one marginal operator. As a consequence, two equations are needed to describe the RG flow in the neighborhood of some critical points. In this case, the general equations describing two-parameter scaling (2PS) can be written using $\beta$ itself as a local coordinate:
@@ V Numerical results
Figure 2: (a) System-size dependence of the fractal dimension $D$ at the center of the spectrum (i.e. at infinite temperature). Dots are the numerical data, obtained by taking the finite-difference derivative of the participation entropy, and solid lines are the interpolation of the data, used for computing the beta function. For any $W$, the curves will eventually reach $D=1$ for sufficiently large system size, as all curves flow towards increasing values of $D$ as the system size grows. At intermediate $L$, the fractal dimension becomes $D>1$, signaling that the eigenfunction support set grows faster than the system size. (b) Beta function at $E=0$, with direction of the RG flow indicated by the arrows. The red dots in the main plot are the zeros of the beta function at $D>1$, and we denote with $D_{B}$ the corresponding value of $D$. In the inset, we report the dependence of $D_{B}$ on $W$: a linear behavior is observed, suggesting the persistence of this effect for large sizes. (c) Pictorial representation of the two possible scenarios for the RG flow at the ergodic fixed point, one-parameter (1PS) or two-parameter (2PS) scaling. The red dashed line is a sketch for the 1PS asymptotic curve.
Numerical data qualitatively agree with the picture coming from perturbation theory [54, 55]. The fractal dimension at the center of the spectrum, $E=0$, is reported as a function of system size in Fig. 2(a), and the corresponding beta function in Fig. 2(b): all curves flow to the ergodic fixed point for any value of $W$, as long as one keeps $W=O(1)$. The same conclusion can be reached from the $r$-parameter data, that we report in the Supplement [73]. Notice that the 2PS cannot be inferred from the flow in the $(r,D)$ plane, as the two parameters are a function of each other [49]. We anticipate that one can observe a transition for $E=0$, provided one rescales the disorder as already pointed out in Refs. [54, 55]. We will explore this transition in the next section.
@@ VI Rescaling the disorder
We plot the beta function of the rescaled $r$-parameter, parametrized by the rescaled disorder $\overline{W}$, in Fig. 3. While bigger system sizes are needed to draw definitive conclusions on the nature of the transition, it appears plausible that it is described by a 2PS theory, with a line of fixed points on the negative $\beta$ axis. The situation appears similar to the Anderson transition on expander graphs [49], and potentially to the random-field XXZ spin chain [57]. The approach to the ergodic fixed point follows instead a 1PS curve, as the beta function curves collapse on one another.
[…]
@@ VII On the mobility edge
Figure 4: (a) Beta function (main) and the corresponding fractal dimension (inset) away from the center of the spectrum, at $E=E_{\mathrm{max}}/4$ with $E_{\mathrm{max}}$ largest energy eigenvalue. The approach to $D=1$ suffers from less severe finite-size effects, and $D$ only slightly exceeds 1. Moreover, at large $W$ the curves $D(L)$ approach $D=0$, signaling the presence of the localized phase. (b) Determination of the critical value $W_{c}$ at energy $E=E_{\mathrm{max}}/4$: $\phi_{A}$ is such that $\beta(\phi_{A})=0$, and it extrapolates to $\phi_{A}=0$ for $W_{c}\simeq 20$. (Inset) At $W=W_{c}$, quantities as the fractal dimension $D$ vanish as $1/L^{2}$, a behavior typical of expander graphs [47, 49].
@@ VIII Discussion
The scaling theory of the localization transition in the QREM has been studied in the present work by reconstructing numerically the beta function of spectral observables. In the “natural” scaling of the disorder variance, which makes the model a mean-field version of MBL models, our analysis confirms the absence of transition at the center of the spectrum and opens two possible scenarios for the approach to the ergodic fixed point: either a one- or a two-parameter scaling. On the other hand, at finite energy density a localization transition is described by a two-parameter scaling theory similarly to the Anderson model on expander graphs. It was noted in Ref. [79] that a different microcanonical phase diagram was proposed, suggesting the existence of a non-ergodic extended phase at finite energy density. We emphasize that our findings are not necessarily in conflict with this conjecture. However, a definitive confirmation of such a phase would require access to larger system sizes, which are essential for a more reliable extrapolation of the $\beta$ function within the ergodic regime.
A transition to a localized phase emerges also at the center of the spectrum if one suitably rescales the variance of the disorder: the transition appears to be described by a two-parameter scaling theory even in this case, with the same critical exponent of the Anderson model on the RRG and possibly of the XXZ spin chain. Our results show that the universality class of the localization transition in the QREM is the same of the Anderson model on any random graph and, remarkably, does not change by rescaling the disorder. This is a consequence of the fact that the RG is blind to the scaling of the microscopic variables.
@@ Acknowledgements.
We acknowledge discussion with C. L. Baldwin, I. V. Gornyi, D. A. Huse, C. R. Laumann, A. D. Mirlin and T. Scoquart, and we thank B. L. Altshuler and V. E. Kravtsov for collaborations on related topics. We are also grateful to one of the Referees for pointing out the agreement between our results and those of Ref. [66]. Numerical simulations were performed using the libraries SciPy [80], PETSc [81] and SLEPc [82] on the MPIPKS HPC cluster. This work was partly funded by the European Union – NextGenerationEU under the project NRRP “National Centre for HPC, Big Data and Quantum Computing (HPC)” CN00000013 (CUP D43C22001240001) [MUR Decree n. 341–15/03/2022] – Cascade Call launched by SPOKE 10 POLIMI: “CQEB” project. The work of JN was funded by the European Union–NextGenerationEU under the project NRRP Project “National Quantum Science and Technology Institute" — NQSTI, Award Number: PE00000023, Concession Decree No. 1564 of 11.10.2022 adopted by the Italian Ministry of Research, CUP J97G22000390007.